OpenStudy (anonymous):

Solve the following inequality : 4/5 x < -1/5 a. x < 1/4 b. x > 1/4 c. x < -1/4 d. x > -1/4

OpenStudy (lalaly):

c

OpenStudy (anonymous):

\[\frac{4}{5}x < -\frac{1}{5}\] \[x < -\frac{1}{4}\]